OpenStudy (anonymous):

Find the distance from the given point to the given line. Estimate your answer to the nearest 2 decimal places.

OpenStudy (anonymous):

There are many points to create a line. Which point on the line do you want to find the distance to?

OpenStudy (anonymous):

The distance formula: \[ d = \sqrt{(x_1-x_2)^2+(y_2-y_1)^2} \] Let \((x_1, y_1) = (x, -2x+7)\) and \((x_2, y_3) = (-6, 4)\) \[ d(x) = \sqrt{[(x)-(-6)]^2+[(-2x+7)-(4)]^2} \] This is the distance with respect to \(x\). Use calculus to minimize this equation. The distance would be the minimum value.

OpenStudy (anonymous):

distance formula is just pythogeoren theriom

OpenStudy (anonymous):

Another way you could do this is to create a line which is perpendicular to \(y=-2x+7\) and goes through the point \((-6, 4)\). Then find the intersection between this perpendicular line and \((-6, 4)\). This intersecting point is going to be the closes point on \(y=-2x+7\) to \((-6, 4)\). So just plug those two points into the distance formula.

OpenStudy (anonymous):

Yet another way to do this is to rotate the line and the point to be parallel to the x axis. Then the line would be a constant \(y=c\) and the point would be relocated to \((x_0, y_0)\). The distance would just be \(|c-y_0|\) in this case.

OpenStudy (anonymous):

plug what?

OpenStudy (anonymous):

Which method are you using?

OpenStudy (anonymous):

idk .. I'm confused here

OpenStudy (anonymous):

could someone please break it down?

OpenStudy (anonymous):

Usually your class should prescribe a method to use... what are you learning in class right now?